(A9). The mean number of errors due to a bug occurring in a minute is 0.0001. What is the
probability that no error will occur in 20 minutes?
1
Expert's answer
2020-06-04T20:36:02-0400

Let X = the random variable denoting the number of errors due to the bug occurring in a minute

The values of X being 0, 1, 2, ..... "\infty"


The mean number of errors due to the bug occurring in a minute = 0.0001


Then, X ~ Poisson("\lambda" = 0.0001) [since, for Poisson distribution parameter = mean]


Then the p.m.f. of X is given by,


P(X = x) = "\begin{cases}
 \frac{e^{-0.0001}.(0.0001)^x}{x!} &\text{for } x=0,1,2,...\infty \\
 0 &\text{otherwise }
\end{cases}"


Now, the probability that no error will occur in 1 minute


= P(X = 0)


= "\frac{e^{-0.0001}.(0.0001)^0}{0!}"


= e-0.0001


"\therefore" The probability that no error will occur in 20 minutes


= "\displaystyle\prod_{i=1}^{20}" (The probability that no error will occur in ith minute)


= [P(X = 0)]20


= [e-0.0001]20


= e-0.002


= 0.998 (rounded to 3 decimal places)


Answer: The probability that no error will occur in 20 minutes is 0.998.